x + y = 16 x-y = 10 solve by substitution.
The answer sheet says (8,-2) I am coming up with something different. Please help!

Answer by addingup(3677)

By substitution:
x + y = 16 subtract y from both sides:
x = 16-y
:
x-y = 10 here substitute for x:
16-y-y = 10
-2y = -6
y = 3 and since 16-y = x then 16-3 = x = 13
-------------------
by addition:
x+y = 16
x-y = 10
2x = 26
x = 13 and since x+y = 16, then 16-13 = y = 3
----------------
As you see, your answer is (13, 3)